# Serializing Java Objects in Plain Code

### v1

@@ 3.4 Research Questions
In order to answer this RQ, we report the number of plain-code serialized objects for each project under the considered workload (#Obj Plain Code Success). We also present the number of objects serialized solely through the structure-based strategy (sb), as well as those serialized using the trace-based strategy (tb). The findings from this RQ illustrate plain-code serialization in action, and highlight the ability of ProDJ in achieving plain-code serialization for real-world Java projects.
RQ2 (Behavior)
@@ 4 Experimental Results
Project RQ1: Objects RQ2: Behavior
@@
Success
sb tb #Tests Pass Fail
@@
PDFBox 100,830 29,385 71,445 489 376 113 4
GraphHopper 40,859 38,325 2,534 258 191 67 3
BroadLeaf 30,273 13,929 16,344 904 403 501 2
Gephi 2,737 1,422 1,315 82 68 14 2
Total 174,699 83,061 91,638 1,733 1,038 695 2
@@ 4.1 RQ1 (Objects)
Columns 2 to 4 of Table 3 present the results for RQ1.First, we note that ProDJ successfully plain-code serializes production objects for all $4$ projects. For instance, ProDJ successfully plain-code serializes $100,830$ objects for PDFBox. Overall, ProDJ successfully plain-code serializes $174,699$ objects that occur at the considered serialization points. Column 3 highlights that nearly 48% of the objects ($83,061$) are plain-code serialized using the structure-based strategy. This means that ProDJ synthesized a minimal reconstruction plan for each of these objects, and instantiated the meta-variables within it, at runtime. Next, from column 4, we see that $91,638$ objects (52%) were serialized via the trace-based strategy. As described in subsection 2.3, this means that ProDJ either found a static field with the same value, or recorded constructor and mutating method calls that can reconstruct the object to its target state.
To illustrate plain-code serialization, 8 shows an excerpt of the serialized version of a GHRequest object in GraphHopper. This serialized object is concise, and we note that the variable names refer to the semantics of the program. The names of the variables, such as points1, correspond to the names of the parameters they are passed as, and are used thanks to source code analysis happening together with plain-code serialization.
@@
[nobreak=true,style=mpdframe,frametitle=Answer to RQ1] ProDJ plain-code successfully serializes thousands of objects ($174,699$) observed at runtime in $4$ real-world Java applications. Both the structure-based and trace-based serialization strategies are useful to translate runtime objects in plain Java.
Listing 8: Excerpt of the plain-code serialized GHRequest object from GraphHopper, requesting a route between two points. Plain-code serialization uses a combination of the structure- and trace-based strategies.
@@
4.2 RQ2 (Behavior)
Listing 9: A test generated by ProDJ for GraphHopper
@@ 4.2 RQ2 (Behavior)
Listing 10: A test generated by ProDJ for PDFBox. A part of the test has been outlined to the helper method createPDColor, that recreates production objects
@@
Per RQ1, the total number of objects plain-code serialized by ProDJ is $174,699$ but we use only a fraction of them in the $1,733$ generated tests. We de-duplicate identical objects, before including them within the generated tests. Furthermore, as trace-based serialization indiscriminately records all instances of the types we monitor, we serialize many objects only for trace-based serialization events. We discard the objects that are not associated with the method under test. This means that we do not use all of the $174,699$ plain-code serialized objects within the $1,733$ generated tests. We present the median number of objects that are deserialized in each each generated test, in the last column of Table 3. For example, $55$ tests generated for PDFBox use $1$ plain-code serialized object, while one generated test uses $122$ deserialized objects, resulting from the reconstruction of a List object and its constituent elements. Overall, the median number of objects in a test generated for PDFBox is $4$.
@@
In 9, we present one test generated by ProDJ for GraphHopper, clearly following the Arrange-Act-Assert pattern. This test verifies the behavior of the alignOrientation method of the AngleCalc class in GraphHopper. ProDJ generates the Java statements from lines $5$ to $7$ at runtime using the structure-based strategy. Note that ProDJ reuses the parameter names of alignOrientation for the local variables in the test on lines $6$ and $7$. The assertion on line $11$ verifies that the double value output from the invocation of alignOrientation on line $9$ matches the value recorded at runtime. We also present one ProDJ-generated test for PDFBox in 10. The method under test, called getComponents defined in the class PDColor, returns a non-null, float array output, based on the value of a field called colorSpace. Note that ProDJ outlines the recreation of the receiver object of type PDColor into a separate createPDColor method. Within this method, ProDJ uses object reconstruction mixing for plain-code serializing PDColor. Specifically, during the pre-execution phase, ProDJ discovers that the associated type COSArray is eligible for trace-based serialization. At runtime, ProDJ records a constructor call event for COSArray (recreated on line $15$), followed by a call to the method addAll (line $22$). The only parameter of addAll is a list, which ProDJ serializes using the structure-based strategy (line $20$). However, the contents of this list are serialized using the trace-based strategy (lines $16$ $18$). ProDJ uses the parameter name objectList of addAll to derive the names of the local variables in the helper method. Note that ProDJ observes the calls to the factory method COSFloat.get at runtime and therefore recreates these calls, instead of invoking the COSBase constructor directly. The String arguments for these COSFloat.get method calls are directly inlined, in contrast with the double arguments on lines $6$ and $7$ in the first test of 9. This is because ProDJ tries to inline primitive values, but only if this would not break up the Arrange-Act-Assert pattern. Next, on line $23$ of 10, a PDColorSpace object is reconstructed using the runtime trace as being an instance of the RGB color space. Finally, the two-argument constructor call on line $24$ reuses the reconstructed objects to recreate the runtime PDColor object. This object is used within the generated test as the receiver for the invocation of getComponents (line $8$). Note also the inline creation of the float array on line $10$ within the assertion, which verifies the equality of the expected and actual float array output. We argue that these plain-code serialization and readability optimization strategies employed by ProDJ result in tests that not only reflect production observations, but are also focused, well-structured, an easy-to-read.
@@
All test cases generated by ProDJ compile correctly and when we execute them, $1,038$ test cases pass. Specifically, for three of the four case studies, a large majority of the generated tests pass. That is a strong property. Given the structure of generated tests, it means that the state of an object after calling a method, i.e., after the ‘Act’ phase of the test, is consistent with the state observed in production (the ‘Assert’ phase). While not being a correctness proof, this is strong consistency evidence. In addition, manual analysis of a random sample of plain-serialized objects in tests, such as the one in 8, confirm that they are meaningful.
@@
[nobreak=true,style=mpdframe,frametitle=Answer to RQ2] Our experiment with using plain-code serialization for testing shows that ProDJ faithfully recreates runtime states. The assertions in this testing context provide soundness evidence.
@@ 4.3 RQ3 (Performance)
The results for RQ3 are presented in Table 4, with the three column groups representing the three measurement categories, corresponding to each of the ProDJ phases discussed in subsection 2.4. We run each experiment three times and keep the lowest measurement from the three runs. This is motivated by the fact that the experimental machine introduces noise in the measurements, which only adds overhead on measurements. Hence, the lowest measurement is the most appropriate measurement of the actual time. Before the application is run, ProDJ analyses the source code. The time to perform this step is presented in the column Pre-execution. This includes parsing the source code, building an AST, and extracting all necessary data for structure- and trace-based serialization. We compute the values by taking the total runtime of this step and dividing it by the amount of types in the project. For example, during the pre-execution phase, ProDJ analyzes $1,883$ types within BroadLeaf, spending $11$ ms per type. On the other hand, ProDJ spends $24$ ms on each of the $590$ types in GraphHopper. This difference is mostly because the analysis overhead per type is less than the parsing overhead. As BroadLeaf is a bigger project, the cost is amortized over a larger number of diverse types.
@@
PDFBox 17 12 4.16 35 105 51
GraphHopper 24 14 5.42 41 137 35
BroadLeaf 11 21 4.38 147 43 38
Gephi 10 18 4.7 85 265 22
@@
The last two columns of Table 4 present the time taken by ProDJ for post-production. This duration includes the time spent offline, analysing the event sequence for objects using trace-based serialization in order to reconstruct them, as well as the time taken taken for readability optimizations, and writing of the post-processed generated tests to disk. The post-execution phase takes under a minute for all projects. Of the four projects, ProDJ spends most time generating tests for PDFBox because the tests contain many mutator calls that are made on objects in PDFBox. Furthermore, the arguments to these calls often consist of structure-based serialized arrays which have multiple entries. As ProDJ needs to build an AST representation of a variable’s type in order to deduplicate objects, this necessitates a lot of string parsing. The time per test is the highest for Gephi, as many tests ProDJ tries to generate contain objects it is unable to fully reconstruct. Consequently, the test has to be discarded, increasing the time per successfully generated test.
@@
5 Discussion
@@
Internal classes Serialization may fail because of JDK-internal classes, such as those loaded by the bootstrap-classloader. When a bootstrap class is instrumented, additional code is inserted and the bootstrap-classloader is used to look up any required ProDJ classes. As the bootstrap-classloader only loads JDK classes, this does not work and the application crashes. It would be possible to include our agent in the bootstrap-classpath, but this sophistication is out of the scope of a research prototype.
@@
Streams Input streams are problematic as they often read from an ephemeral source, such as a network stream, which might no longer exist when the object is reconstructed. A simple case illustrating this problem is file access or byte array input streams, as can be seen with PDFBox. In this case, ProDJ fails to serialize java.io.ByteArrayInputStreams and org.apache.fontbox.afm.CharMetric.
@@
Runtime code generation Runtime annotation processing cannot be handled. In Java, annotations are used for dynamical analysis, and for generating implementations or proxy classes on-the-fly. These classes do not exist on disk, and can therefore not be plain-code serialized.
@@
Bounded depth In our experiments, ProDJ limits the length of serialized arrays to $25$, in order to have readable test cases. Consequently, when a test case asserts behavior that requires more than $25$ elements, the assertion fails. For example, $37$ test cases generated for PDFBox fail for this reason.
@@
Visibility The next challenge relates to visibility and encapsulation, as some fields and methods may not be visible in source code. For example, the trace-based reconstruction strategy can capture calls to package-private methods, which cannot be replayed in the test. We observed such a limitation for $22$ tests of PDFBox. To address this limitation, ProDJ would need to determine the highest accessible method creating objects of a given type.
@@
Environment state The last major challenge for serialization relates to parts of the state that are related to the execution environment. For example, $11$ tests in PDFBox fail because they try to read a certain file, which does not exist anymore during test execution. ProDJ only recreates states that are local to serialized objects, not the complete system state.

### v3

@@ 3.4 Research Questions
In order to answer this RQ, we report the number of plain-code serialized objects for each project under the considered workload (#Obj Plain Code). We also present the number of objects serialized solely through the structure-based strategy (sb), as well as those serialized using the trace-based strategy (tb). The findings from this RQ illustrate plain-code serialization in action, and highlight the ability of ProDJ in achieving plain-code serialization for real-world Java projects.
RQ2 (Accuracy)
@@
RQ4 (User Study)
What is the opinion of developers on tests containing plain-code serialized objects?
With this RQ, our goal is to assess the opinion of developers on the representation of serialized objects as plain code, relative to other serialized textual representations, specifically XML and JSON. In order to do this, we first select a set of two generated tests for each of the four projects under study, which contain objects serialized at runtime. Next, we invite 17 developers in our professional network over email, to participate in an online survey. As part of the survey, we present each developer a set of five random pairs of generated tests. Each pair contains two variants of the same test, contrasting XML or JSON with ProDJ-produced plain-code serialized objects. This means that one test in the pair contains either XML or JSON, while the other contains ProDJ-produced plain-code serialized objects. We invite each participant to mark the test in each pair that is more readable, according to their preference. For each pair of tests that receives two responses, we compute the Cohen’s kappa coefficient for agreement. For pairs that receive more than two responses, we report the Fleiss’ kappa [22]. We analyze the responses from this survey and present our findings in subsection 4.4. Our replication package contains all details of this survey.
@@ 4 Experimental Results
Project RQ1: Objects RQ2: Accuracy
@@
sb tb #Tests Pass
@@
PDFBox 100,830 29,385 71,445 489 376 (82%) 4
GraphHopper 40,859 38,325 2,534 258 191 (74%) 3
BroadLeaf 30,273 13,929 16,344 904 403 (45%) 2
Gephi 2,737 1,422 1,315 82 68 (82%) 2
Total 174,699 83,061 91,638 1,733 1,038 2
@@ 4.1 RQ1 (Objects)
Columns 2 to 4 of Table 3 present the results for RQ1. First, we note that ProDJ successfully plain-code serializes production objects for all $4$ projects. For instance, ProDJ plain-code serializes $100,830$ objects for PDFBox. Overall, ProDJ plain-code serializes $174,699$ objects that occur at the considered serialization points. Column 3 highlights that nearly 48% of the objects ($83,061$) are plain-code serialized using the structure-based strategy. This means that ProDJ synthesized a minimal reconstruction plan for each of these objects, and instantiated the meta-variables within it, at runtime. Next, from column 4, we see that $91,638$ objects (52%) were serialized via the trace-based strategy. As described in subsection 2.3, this means that ProDJ either found a static field with the same value, or recorded constructor and mutating method calls that can reconstruct the object to its target state.
To illustrate plain-code serialization, 9 shows an excerpt of the serialized version of a GHRequest object in GraphHopper. This serialized object is concise, and we note that the variable names refer to the semantics of the program. The names of the variables, such as points1, correspond to the names of the parameters they are passed as, and are used thanks to source code analysis happening together with plain-code serialization.
@@
[nobreak=true,style=mpdframe,frametitle=Answer to RQ1] ProDJ successfully plain-code serializes thousands of objects ($174,699$) observed at runtime in $4$ real-world Java applications. Both the structure-based and trace-based serialization strategies are useful to translate runtime objects in plain Java.
Listing 9: Excerpt of the plain-code serialized GHRequest object from GraphHopper, requesting a route between two points. Plain-code serialization uses a combination of the structure- and trace-based strategies.
@@
4.2 RQ2 (Accuracy)
Listing 10: A test generated by ProDJ for GraphHopper
@@ 4.2 RQ2 (Accuracy)
Listing 11: A test generated by ProDJ for PDFBox. A part of the test has been outlined to the helper method createPDColor, that recreates production objects
@@
Per RQ1, the total number of objects plain-code serialized by ProDJ is $174,699$ but we use only a fraction of them in the $1,733$ generated tests. We de-duplicate identical objects, before including them within the generated tests. We consider an object to be identical to another if the generated statements are the same, modulo identifier names, or refer to the same object (==) reconstructed using trace-based serialization. Furthermore, as trace-based serialization indiscriminately records all instances of the types we monitor, we serialize many objects only for trace-based serialization events. For example, we might record multiple GHRequest instances, but only one is actually used in a method under test. We then discard all other instances that are not associated with it. This means that we do not use all of the $174,699$ plain-code serialized objects within the $1,733$ generated tests. We present the median number of objects that are deserialized in each each generated test, in the last column of Table 3. For example, $55$ tests generated for PDFBox use $1$ plain-code serialized object, while one generated test uses $122$ deserialized objects, resulting from the reconstruction of a List object and its constituent elements. Overall, the median number of objects in a test generated for PDFBox is $4$.
@@
In 10, we present one test generated by ProDJ for GraphHopper, clearly following the Arrange-Act-Assert pattern. This test verifies the behavior of the alignOrientation method of the AngleCalc class in GraphHopper. ProDJ generates the Java statements from lines $5$ to $7$ at runtime using the structure-based strategy. Note that ProDJ reuses the parameter names of alignOrientation for the local variables in the test on lines $6$ and $7$. The assertion on line $11$ verifies that the double value output from the invocation of alignOrientation on line $9$ matches the value recorded at runtime. We also present one ProDJ-generated test for PDFBox in 11. The method under test, called getComponents defined in the class PDColor, returns a non-null, float array output, based on the value of a field called colorSpace. Note that ProDJ outlines the recreation of the receiver object of type PDColor into a separate createPDColor method. Within this method, ProDJ uses object reconstruction mixing for plain-code serializing PDColor. Specifically, during the pre-execution phase, ProDJ discovers that the associated type COSArray is eligible for trace-based serialization. At runtime, ProDJ records a constructor call event for COSArray (recreated on line $15$), followed by a call to the method addAll (line $22$). The only parameter of addAll is a list, which ProDJ serializes using the structure-based strategy (line $20$). However, the contents of this list are serialized using the trace-based strategy (lines $16$ $18$). ProDJ uses the parameter name objectList of addAll to derive the names of the local variables in the helper method. Note that ProDJ observes the calls to the factory method COSFloat.get at runtime and therefore recreates these calls, instead of invoking the COSBase constructor directly. The String arguments for these COSFloat.get method calls are directly inlined, in contrast with the double arguments on lines $6$ and $7$ in the first test of 10. This is because ProDJ tries to inline primitive values, but only if this would not break up the Arrange-Act-Assert pattern. Next, on line $23$ of 11, a PDColorSpace object is reconstructed using the runtime trace as being an instance of the RGB color space. Finally, the two-argument constructor call on line $24$ reuses the reconstructed objects to recreate the runtime PDColor object. This object is used within the generated test as the receiver for the invocation of getComponents (line $8$). Note also the inline creation of the float array on line $10$ within the assertion, which verifies the equality of the expected and actual float array output. We argue that these plain-code serialization and readability optimization strategies employed by ProDJ result in tests that not only reflect production observations, but are also focused, well-structured, an easy-to-read.
@@
All test cases generated by ProDJ compile correctly and when we execute them, $1,038$ test cases pass, which is a success rate of $59.9$%. Specifically, for three of the four case studies, a large majority of the generated tests pass. For PDFBox, $82$% of all tests generated pass, the highest among the four projects. That is a strong property. Given the structure of generated tests, it means that the state of an object after calling a method, i.e., after the ‘Act’ phase of the test, is consistent with the state observed in production (the ‘Assert’ phase). While not being a correctness proof, this is strong consistency evidence. In addition, manual analysis of a random sample of plain-serialized objects in tests, such as the one in 9, confirm that they are meaningful.
@@
[nobreak=true,style=mpdframe,frametitle=Answer to RQ2] Our experiment with using plain-code serialization for testing shows that ProDJ faithfully recreates runtime states for all objects in nearly 60% of the generated tests. The assertions in this testing context provide soundness evidence.
@@ 4.3 RQ3 (Performance)
The results for RQ3 are presented in Table 4, with the three column groups representing the three measurement categories, corresponding to each of the ProDJ phases discussed in subsection 2.4. We run each experiment three times and present the median measurement from the three runs. Before the application is run, ProDJ analyses the source code. The time to perform this step is presented in the column Pre-execution. This includes parsing the source code, building an AST, and extracting all necessary data for structure- and trace-based serialization. We compute the values by taking the total runtime of this step and dividing it by the amount of types in the project. For example, during the pre-execution phase, ProDJ analyzes $1,883$ types within BroadLeaf, spending $11$ ms per type. On the other hand, ProDJ spends $27$ ms on each of the $590$ types in GraphHopper. This difference is mostly because the analysis overhead per type is less than the parsing overhead. As BroadLeaf is a bigger project, the cost is amortized over a larger number of diverse types.
@@
PDFBox 18 14 0.14 35 105 59
GraphHopper 27 16 0.40 41 65 40
BroadLeaf 11 23 0.78 157 38 40
Gephi 11 21 0.51 159 145 29
@@
The last two columns of Table 4 present the time taken by ProDJ for post-production. This duration includes the time spent offline, analysing the event sequence for objects using trace-based serialization in order to reconstruct them, as well as the time taken taken for readability optimizations, and writing of the post-processed generated tests to disk. The post-execution phase takes under a minute for all projects. Of the four projects, ProDJ spends most time generating tests for PDFBox because the tests contain many mutator calls that are made on objects in PDFBox. Furthermore, the arguments to these calls often consist of structure-based serialized arrays which have multiple entries. As ProDJ needs to build an AST representation of a variable’s type in order to de-duplicate objects, this necessitates a lot of string parsing. The time per test is the highest for Gephi, as many tests ProDJ tries to generate contain objects it is unable to fully reconstruct. Consequently, the test has to be discarded, increasing the time per successfully generated test.
@@
4.4 RQ4 (User Study)
We conduct a user study per the protocol defined in subsection 3.4, based on an online survey with developers. The goal of the survey is to assess the opinion of developers regarding the readability of tests containing plain-code serialized objects. This survey received 17 responses by developers with 20 years of median experience with programming, and 13 years programming with Java. We prepare 16 pairs of tests in total, presenting five random pairs to each respondent. We select two tests for each project which have less than 30 lines of Java code, including outlined utility methods. For the JSON and XML variants of these tests, we collapsed identical lines in the JSON and XML code. The maximum number of lines in the tests with JSON and XML tests was 1,797 and 1,787, respectively. Each respondent was shown five randomly selected pairs. One test within the pair contains plain-code serialized objects, while the other was the same test, but with either JSON- or XML-serialized objects. We randomized the order in which the plain-code and JSON or XML tests were presented. The respondents were requested to mark which test in each pair they preferred, with respect to its readability. Moreover, 11 of the 17 respondents shared additional qualitative comments about serialization.
Project Test pair Votes C or F kappa (agreement)
Plain-code JSON/XML
PDFBox PDF1–JSON 8 0 F 1.0 (perfect)
PDF1–XML 3 0 F 1.0 (perfect)
PDF2–JSON 4 2 F -0.07 (poor)
PDF2–XML 6 1 F 0.43 (moderate)
GraphHopper GH1–JSON 7 1 F 0.5 (moderate)
GH1–XML 6 1 F 0.43 (moderate)
GH2–JSON 2 0 C 1.0 (perfect)
GH2–XML 6 0 F 1.0 (perfect)
BroadLeaf BL1–JSON 4 1 F 0.2 (slight)
BL1–XML 9 0 F 1.0 (perfect)
BL2–JSON 5 0 F 1.0 (perfect)
BL2–XML 2 0 C 1.0 (perfect)
Gephi GEP1–JSON 4 0 F 1.0 (perfect)
GEP1–XML 6 0 F 1.0 (perfect)
GEP2–JSON 3 0 F 1.0 (perfect)
GEP2–XML 4 0 F 1.0 (perfect)
Total 16 79 6
Table 5: Summary of our user study. We select 2 tests from each of the four projects for the study. The second column lists the 16 test pairs we prepare using these tests. Each pair contains one test with plain-code serialized objects and an equivalent test with either JSON or XML representations of the same serialized objects. We present 5 random test pairs to each of the 17 participants of the study, who vote for tests with plain-code (column 3) or JSON/XML (column 4) serialized objects, based on their readability. The last column presents the Cohen’s (C) or Fleiss’ (F) kappas of agreement between the participants.
Table 5 summarizes the responses from the survey. We find that through 79 of the total 85 votes (92.9%), the developers showed a preference for the tests containing the plain-code serialized objects. In only 6 cases, they preferred the versions with XML or JSON representations. For 11 of the 16 test pairs, the preference for plain-code serialized objects was unanimous. In those cases, the agreement measure (Cohen’s and Fleiss’ kappas) is a perfect 1.0. In the 5 remaining cases, the agreement varies from poor to moderate.
Qualitative statements by the developers confirmed the preference, as the follows:
[…]
[4pt]       I [prefered] the shorter versions because the longer ones contained serialization ”noise” that did not seem relevant.
[…]
[4pt]       Intuitively, the [plain-code version] is often more readable. However, legacy code may be hard to test as classes could have many dependencies that are hard to construct. In those cases, the readability gap between object construction and deserialization could be narrowed.
A single developer voted for the XML or JSON variants of all presented pairs, hence they are responsible for 5 of the 6 votes for JSON/XML, owing to the higher likelihood of capturing unintentional breakages:
[…]
[4pt]       I prefer the unchanging nature of JSON and XML, which do not get refactored automatically in sync with the application code, when the classes under test change.
The test pair PDF2–JSON has the lowest agreement score. This is due to the aforementioned developer who prefers JSON/XML, and another developer who misunderstood the meaning of the test case, per their comment. Note that the latter respondent ranked plain-code tests higher in the other four pairs they were shown.
One developer correctly noted that the tests would be organized differently in practice:
[…]
[4pt]       I believe I would store the long text-snippets as variables in another file, instead of having them in the test.
Finally, a respondent balances the requirements of serialization performance versus readability.:
[…]
[4pt]       When sending data back and forth over the wire, more efficient formats may be better, like Protobuf […]. Overall, it really depends on the task at hand.
This is valid point Our technique is entirely optimized for readability, and we do recommend binary serialization for performance-sensitive use cases.
{mdframed}
[nobreak=true,style=mpdframe,frametitle=Answer to RQ4] Our user study demonstrates, through a sound protocol, that developers rate tests containing plain-code serialized objects higher with respect to their readability, in comparison with equivalent tests with XML- or JSON-serialized objects.
5 Limitations and Threats
@@
Serialization may fail because of JDK-internal classes, such as those loaded by the bootstrap-classloader. When a bootstrap class is instrumented, additional code is inserted and the bootstrap-classloader is used to look up any required ProDJ classes. As the bootstrap-classloader only loads JDK classes, this does not work and the application crashes. It would be possible to include our agent in the bootstrap-classpath, but this sophistication is out of the scope of a research prototype.
Streams
@@
Input streams are problematic as they often read from an ephemeral source, such as a network stream, which might no longer exist when the object is reconstructed. A simple case illustrating this problem is file access or byte array input streams, as can be seen with PDFBox. In this case, ProDJ fails to serialize java.io.ByteArrayInputStreams and org.apache.fontbox.afm.CharMetric.
Runtime code generation
@@
Runtime annotation processing is not handled by ProDJ. In Java, annotations are used for dynamic analysis, and for generating implementations or proxy classes on-the-fly. These classes do not exist on disk, and can therefore not be plain-code serialized. Likewise, for projects such as BroadLeaf that heavily rely on dependency injection through runtime code generation, ProDJ can not faithfully recreate all objects.
Bounded depth
@@
ProDJ uses a bounded serialization approach to limit the size of objects and consequently the size of generated tests. In our experiments, we limit the length of serialized arrays to $25$, in order to have readable test cases. Consequently, when a test case asserts behavior that requires more than $25$ elements, the assertion fails. For example, $37$ test cases generated for PDFBox fail for this reason.
Visibility
@@
The next challenge relates to visibility and encapsulation, as some serialized fields and methods may not be visible in source code. For example, the trace-based reconstruction strategy can capture calls to package-private methods, which cannot be replayed in the test. We observed such a limitation for $22$ tests of PDFBox. To address this limitation, ProDJ would need to determine and only use accessible method creating objects of a given type. A similar limitation occurs when fields are set through reflection or native code. Such reconstructions are not supported by ProDJ.
Security Implications
Untrusted object deserialization poses a significant security risk that has been increasingly prevalent in recent years [23]. This holds for plain-code representations of runtime objects observed in the field. An attacker could potentially craft a special object, on which they execute a specific malicious sequence of method calls on the testing or CI machine, which is then captured and replayed by ProDJ. Developers should therefore be mindful of the risks associated with deserialization of production objects. On the positive side, we note that plain code is easier to audit for security than alternative serialization formats.
State
@@
The last major challenge for serialization relates to identify all parts of the state that are related to the execution. For example, $11$ tests in PDFBox fail because they try to read a certain file, which does not exist anymore during test execution. ProDJ only recreates states that are local to serialized objects, not the complete system state. This is also a potential challenge at reconstruction time, reconstructing a serialized object may incur side effects that we do not handle, such as writing on disk.
Threats to Validity
Threats to the internal validity of our findings may arise from the tools that we use to implement ProDJ. We mitigate this threat by relying on open-source state-of-the-art tools to handle instrumentation and code transformation.
External threats to the validity of our findings are related to the artifacts we use for our experiments. We demonstrate the feasibility of plain-code serialization of runtime objects through four large, real-world Java projects exercised with representative field workloads. We envision that our approach will also be applicable to other Java projects.
We note that the benefits of plain-code serialization may not generalize to all use cases of serialization. For example, inter-process communication involves the transmission of serialized objects that are not necessarily required to be human-readable. Serializing these objects in plain code is also likely to be impractical due to performance costs. On the other hand, where readability is in fact the use case, we believe that our strategy can be adapted to handle plain-code serialization in other programming languages. The key requirement is that their runtime should support introspection, which is the case for most modern languages, including JavaScript and Python.